April 11, 2014

19:30 Cloudy: Just done dark calibration and opening the telescope now.

19:38 Telluric acquisition HIP 035050, seeing ~1.4. Clouds

20:05 Lost 12 magnitude guide star

20:10 Acquisition HIP 61534

20:13 MOS error with mask M1532

21:25 Switching to MODS

21:42 Waiting for clouds to clear...

23:38 Dave was able to move the LUCI mask into a safe position, waiting for Peter to take a look at it to fix remaining problems and decide its future use...

04:57 Never reopened, clouds were pretty thick throughout the night.

-- %USERSIG{GregoryLeeWalth - 2014-04-12}%

Comments

 
Topic revision: r5 - 14 Apr 2014, GregoryLeeWalth
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback